jigglypuff314 (jigglypuff314):

\[If~~ f(x) =\int\limits_{0}^{x}(t ^{2}-3t+2) dt\]find f'(x) I got \[f'(x) = \frac{ 1 }{ 3 }x ^{3} - \frac{ 3 }{ 2 }x ^{2}+2x\]would that be correct? or am I totally wrong? XD

zepdrix (zepdrix):

No silly billy that's not right! D: You correctly calculated \(\Large\bf\sf f(x)\) Now take its derivative.

jigglypuff314 (jigglypuff314):

so x^2 - 3x + 2 ?

zepdrix (zepdrix):

Yay good job \c:/

jigglypuff314 (jigglypuff314):

Awesome :D Thanks!

zepdrix (zepdrix):

You could have also applied the `Fundamental Theorem of Calculus: Part 1` to this problem:\[\Large\bf\sf \frac{d}{dx}\int\limits_0^x f(t)\;dt \quad=\quad f(x)\]

zepdrix (zepdrix):

But going through it the long way is good, you get familiar with the process c:

jigglypuff314 (jigglypuff314):

coolios :) I have no idea what I'm doing XD my teacher says there's a short cut? :3

zepdrix (zepdrix):

Yah that's the shortcut :U

zepdrix (zepdrix):

Example Problem:

zepdrix (zepdrix):

\[\Large\bf\sf f(x)\quad=\quad \int\limits_0^x \frac{e^{\arctan(\sqrt[3]{t^2-1})}}{3t^2+7t-1}dt\] Then the derivative is simply,\[\Large\bf\sf f'(x)\quad=\quad \frac{e^{\arctan(\sqrt[3]{x^2-1})}}{3x^2+7x-1}\]No need to do all the fancy work of trying to find an anti-derivative and then differentiate it again.

jigglypuff314 (jigglypuff314):

ohhh so you just replace t with x ? what if it was 0 to x^2 ?

zepdrix (zepdrix):

Ohh good question. Then when you go to differentiate, you have to apply the chain rule.\[\Large\bf\sf \frac{d}{dx}\int\limits\limits_0^{x^2}f(t)\;dt\quad=\quad \frac{d}{dx}F(t)|_0^{x^2}\] \[\Large\bf\sf =\frac{d}{dx}\left[F(x^2)-F(0)\right]\]Differentiationg finally,\[\Large\bf\sf =f(x^2)\frac{d}{dx}x^2-0\quad=\quad f(x^2)\cdot 2x\]

zepdrix (zepdrix):

So as before, we simply stick x^2 in for all of the t's. But now we also get an extra factor of 2x because of the chain rule.

jigglypuff314 (jigglypuff314):

mmm then how would trig work? cuz I was given \[f(x) = \int\limits_{0}^{x ^{2}}(t + \cos 3t)dt\]

zepdrix (zepdrix):

Our integrand is a function, \[\Large\bf\sf f(\color{royalblue}{t})\quad=\quad \color{royalblue}{t}+\cos3\color{royalblue}{t}\]We take it's anti-derivative, giving us some function we'll call,\[\Large\bf\sf F(\color{royalblue}{t})\]Plug in the boundaries, and take our derivative,\[\Large\bf\sf =\quad f(\color{royalblue}{x^2})\cdot 2x\quad=\quad (\color{royalblue}{x^2}+\cos3\color{royalblue}{x^2})\cdot 2x\]

zepdrix (zepdrix):

It's important to understand what is going on. But for the sake of simplification you can think of it as, ~I get the same thing back, but with `the upper bound` inside the function instead of t, ~and i have to multiply by the derivative of the upper bound.

jigglypuff314 (jigglypuff314):

omg that short cut is so cool :D if you don't mind helping me with one more? :3 \[\huge f(x) = \int\limits_{-1}^{xsinx}\sqrt{t ^{2}+1}dt\]

zepdrix (zepdrix):

Do you understand why the lower boundary is giving us zero? (It will happen in this problem as well)

jigglypuff314 (jigglypuff314):

mmm before when lower bound was 0 it wasn't involved cuz it was stuff times 0 which equals 0 right? but now we has a -1 as lower bound so would it be like F(xsinx) - F(-1) ? (has no idea what that notation even means XD)

zepdrix (zepdrix):

So like.. think of a simple example really quick, Because the notation is important. :) \[\Large\bf\sf \int\limits_a^x f(t)\;dt\]We have some stuff inside that we're integrating. We're generalizing by calling it a function. We use big f to denote the anti-derivative. We might not actually be able to find the anti-derivative but that doesn't matter. We just call it F(t). The process is, ~Anti-derivative ~plug bounds in ~take derivative If you integrate something, then differentiate it, you should expect to get back what you started with, yes? It's like you're doing something, then you're undoing it. But with this process we do something between the doing and undoing. We change the argument of our function.

zepdrix (zepdrix):

So when you integrate you get some anti-derivative function,\[\Large\bf\sf F(t)\]Then evaluating it at the upper and lower bounds,\[\Large\bf\sf F(x)-F(a)\]Now when we take a derivative (with respect to x), the second piece that we have there....... is just a constant.

zepdrix (zepdrix):

It's some function (that we're pretending we found), and we're evaluating it at a constant.

zepdrix (zepdrix):

The derivative of a constant is zero, right?

jigglypuff314 (jigglypuff314):

oh right yea :) so it would still be the same idea? what would happen to the -1?

zepdrix (zepdrix):

We'll call our integrand \(\Large\bf\sf f(t)\). So let's assume that some anti-derivative exists. Integrating gives us,\[\Large\bf\sf F(t)\]Evaluating it at the upper and lower bounds gives us,\[\Large\bf\sf F(x \sin x)-F(-1)\]as you said. Now we need to take a derivative.

zepdrix (zepdrix):

If I have the function,\[\Large\bf\sf G(x)=x^2\]And evaluate it at -1,\[\Large\bf\sf G(-1)=1\]What is the derivative of G(-1)?

jigglypuff314 (jigglypuff314):

zero?

zepdrix (zepdrix):

Yes, you simply took the derivative of 1, correct?

jigglypuff314 (jigglypuff314):

yeah :D

zepdrix (zepdrix):

Whenever we take the derivative of a function which has a `constant` plugged in, then it just gives us zero.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{d}{dx}\left[F(x \sin x)-\color{orangered}{F(-1)}\right]\quad=\quad f(x \sin x)\frac{d}{dx}(x \sin x)-\color{orangered}{0}\]

zepdrix (zepdrix):

We get back the same thing we started with, but with (x sin x) plugged in for all the t's. And we also have to multiply by the derivative of (x sin x).

jigglypuff314 (jigglypuff314):

which is xcosx + sinx ?

zepdrix (zepdrix):

Mmmm yah looks good.

zepdrix (zepdrix):

\[\Large\bf\sf\quad=\quad \sqrt{(x \sin x)^{2}+1}\left[x \cos x + \sin x\right]\]

jigglypuff314 (jigglypuff314):

should I simplify or is than an acceptable answer? :3

zepdrix (zepdrix):

That's a good answer. Don't mess with it :)

jigglypuff314 (jigglypuff314):

okay thank you! :D I'll try to do the rest myself, and if I still don't understand I'll ask :) Thank you sooo much! <3

zepdrix (zepdrix):

np ^^ You might run into another form,\[\Large\bf\sf \frac{d}{dx}\int\limits_x^0f(t)\;dt\]Something like that perhaps, where the variable is in the lower limit. You'll want to get it into the upper limit before you can apply this shortcut.

jigglypuff314 (jigglypuff314):

right XD (looking at one now) it'll just be a negative in front right? :)

zepdrix (zepdrix):

Yes good! That's how you switch the bounds :)

jigglypuff314 (jigglypuff314):

so \[f(x) = \int\limits_{x^2}^{4} sint ~dt\]becomes \[f'(x) = -2xsin(x^2 +1)\]

jigglypuff314 (jigglypuff314):

whoops lower bound x^2 + 1

zepdrix (zepdrix):

Oh ok, ya looks correct c:

jigglypuff314 (jigglypuff314):

awesome :) and just to check if I got the concept down... \[\large f(x) = \int\limits_{3-x}^{2x}\sec^5(t) dt\]would be \[\large f'(x) = 2\sec^52x + \sec^5(3-x)\]

zepdrix (zepdrix):

So normally we would `subtract` the lower bound. But it looks like you got an extra negative from the chain rule, so you just combined them. Yessss good job!

jigglypuff314 (jigglypuff314):

YAYYYY!!!! :D *happy dance*

zepdrix (zepdrix):

hah XD